OpenStudy (anonymous):

if1/x,1/y,1/z are in A.P then show that xy.zx,yz are in A.P pls show me the steps of the sum

OpenStudy (anonymous):

$$1/y-1/x=1/z-1/y\\xz-yz=xy-xz$$